# Automated training of neural-network interatomic potentials

### v1

@@ I.3 Validation
Figure 8 presents the phonon dispersion and density of states for graphene, graphite, and diamond. The comparison between NNIP predictions and density-functional perturbation theory (DFPT) [45] calculations demonstrates good agreement across all three carbon allotropes, confirming the potential’s ability to accurately capture vibrational properties. The observed small imaginary ZA phonons near $\Gamma$ are a well-known numerical issue for 2D materials, which can be solved by adopiting prohibitively tight parameters and in particular very high plane-wave cutoffs [46]. Notably, despite being trained on DFT data exhibiting this, the NNIP does not display such unphysical behavior.
II Discussion
@@
The SOREP-based dimensionality reduction (tSNE) has shown that subsequent MD-based active learning steps explore the PES through a dual mechanism: dense sampling of already-known regions and, simultaneously, exploration of entirely new basins. A compelling example is the spontaneous formation of carbon nanotubes during the MD simulations in the active learning cycle: carbon nanotubes were absent from the original dataset but become then automatically incorporated by AiiDA-TrainsPot into subsequent training iterations, suggesting the ability of the automated workflow to find novel and quite different metastable or stable structures without prior knowledge.
@@
It would be interesting to investigate other exploration strategies beyond NPT and NVT MD, such introducing metadynamics [47] by interfacing AiiDA-Trainspot to PLUMED [48, 49, 50]. More in general, we hope that the AiiDA’s plugin system and the modular structure of AiiDA-TrainsPot will encourage and facilitate future upgrades, as well as the integration of new tools. An example would be supporting multiple NNIP engines (beyond MACE) and electronic structure codes (beyond Quantum ESPRESSO), in the spirit of previous efforts on code-agnostic common workflows for EOS and dissociation curves [51]. Powerful upgrades would be enabled by interfacing AiiDA-TrainsPot with existing specialized AiiDA workflows, for instance using DFT+U calculations where the Hubbard U can be automatically calculated for each configuration either with DFPT [52, 53, 54] through the AiiDA-Hubbard workflow [55] or even more efficiently through ML methods [56].
@@
While AiiDA-TrainsPot can operate autonomously for general-purpose NNIP development, domain experts retain full flexibility to incorporate their physical and chemical intuition in the automation strategy. The modular architecture is designed to enable full customization of all key components: initial structure selection, dataset augmentation parameters, MD simulation conditions, and computational settings for integrated codes—Quantum ESPRESSO, MACE, and LAMMPS. Such level of customization allows users to balance the power of automation with the flexibility that is needed to support a wide range of applications, which require tailoring the active learning process to specific research objectives. Indeed, while AiiDA-TrainsPot automates the entire process—traditionally long, tedious and prone to human errors—of developing NNIPs, optimal results still benefit from careful consideration of the system of interest. In other words, the selection of initial structures, augmentation strategies, and MD conditions, can—and often should—be tailored to reflect the target application and desired properties: AiiDA-TrainsPot makes that effort straightforward. On top of that, the enforcement of standardized protocols (either already established, e.g., SSSP pseudopotentials [18] or “fast”/“stringent” QE protocols [21], or introduced in this work) contribute to precision, reproducibility and seamless integration with future efforts in training larger models.
@@
III Methods
All calculations were performed using AiiDA-TrainsPot (version v1.0.0), Quantum ESPRESSO [57, 58, 59] v7.3.1, MACE [5] v0.3.12, and LAMMPS [60] v8Feb2023. The initial dataset included 48 structures, primarily sourced from the Materials Cloud 2D and 3D databases [46, 61], with additional nanoribbons and fullerenes from the Drautz dataset [22].
@@ III Methods
IV Data availability
The training datasets and trained models are available on the Materials Cloud Archive [63].
V Code availability
AiiDA-TrainsPot is available on GitHub at https://github.com/aiida-trieste-developers/aiida-trains-pot.
VI Acknowledgments
@@
VII Author contributions
VIII Competing Interests
The authors declare no competing interests.

### v2

@@ II.3 Validation
We present in Fig. 8 the phonon dispersion and density of states for graphene, graphite, and diamond. The comparison between NNIP predictions and density-functional perturbation theory (DFPT) [4] calculations demonstrates good agreement across all three carbon allotropes, confirming the potential’s ability to accurately capture vibrational properties. The observed small imaginary ZA phonons near $\Gamma$ are a well-known numerical issue for 2D materials, which can be solved by adopting prohibitively tight parameters and in particular very high plane-wave cutoffs [43]. Notably, despite being trained on DFT data exhibiting this, the NNIP does not display such unphysical behavior.
Figure 9: Radial distribution functions $g(r)$ of amorphous carbon structures with a density of 3.5 g/cm3, generated using a melt-and-quench protocol and analyzed with our five NNIPs (thin brown lines) and their committee average (thick black line). The results demonstrate that our potentials accurately reproduce the AIMD, including the positions of the first and second peaks, and the characteristic nonzero minimum at $\sim 1.9$ Å.
In order to further assess the transferability of our NNIP, we consider amorphous carbon, which features a disordered network of mixed sp2 and sp3 bonds that was not explicitly included in the rattled and strained crystalline configurations included in the training set. We compute the radial distribution function (RDF), $g(r)$, which characterizes short-range order in disordered systems, on independent amorphous configurations at a density of 3.5 g/cm3 that have been obtained using the melt-and-quench procedure described in Ref. [60]. Figure 9 compares the RDFs for the committee of potentials and the committee average with reference ab initio molecular dynamics (AIMD) data from Ref. [22]. The NNIPs closely reproduce the AIMD reference, accurately capturing the position and height of the first and second peaks, confirming the ability of our potentials to reliably describe the local structure of amorphous carbon.
Figure 10: Parity plots for energy, forces, and stress on 98% of the SACADA dataset, comparing three models: (i) our MACE-based potentials generated in run B (orange), (ii) the universal PET-MAD potential (blue), and (iii) the PET-MAD potential fine-tuned using the carbon dataset produced in run B (green). The dashed line indicates perfect agreement with the reference DFT data.
As a further transferability test, we evaluate one NNIP on the SACADA dataset, which contains 1635 carbon allotropes collected from the scientific literature [30]. Out of the 1635, we exclude 13 structures due to unconverged DFT labeling. Then, we discard 19 structures where at least one of the potentials yields a committee-based disagreement larger than the following loose thresholds: 0.2 eV for energy, 50 eV/Å for forces, and 1 eV/Å3 for the stress tensor. The remaining 1603 structures (98% of the SACADA dataset) represent the overlap set between converged DFT simulations and successful NNIPs predictions, as required for a meaningful comparison. Here, we also evaluate the universal PET-MAD potential (v1.0.2) [42], both in its original form and after fine-tuning on the carbon dataset generated in run B. Since run B was obtained with PBE functionals, while PET-MAD was trained with PBEsol [48], we relabel the SACADA dataset with PBEsol for evaluating the accuracy of PET-MAD potentials. In passing, we note how the modular structure of AiiDA-TrainsPot permits a seamless exchange of training engines (MACE or Metatrain) and architectures, the reuse of existing labeled datasets, and uniform evaluation procedures across models. We report the parity plots and RMSE in Fig. 10: the MACE potential trained from scratch with AiiDA-TrainsPot on 1795 structures performs similarly to the PET-MAD foundation model, with comparable error on forces and stress tensors while doing slightly worse on energies (89 vs. 54 meV). Notably, fine-tuning PET-MAD on the PBEsol-relabeled run B datasets lowers the error on all metrics (energies, forcess and stress) both compared to the MACE potential trained from scratch and to PET-MAD.
Phase stability in monolayer $\mathrm{W_{x}Mo_{1-x}Te_{2}}$ alloys
As a further validation, we apply AiiDA-TrainsPot to a more challenging problem: the relative phase stability of $\mathrm{W_{x}Mo_{1-x}Te_{2}}$ monolayer alloys. This system is well known to exhibit a composition-dependent transition between the H phase and $\mathrm{T^{\prime}}$ phase, which can be driven by temperature, strain, doping, or alloying [73]. In this benchmark, we specifically assess the ability of different models to reproduce the evolution of the formation-energy difference between the two phases as a function of alloy composition.
For each W fraction $x$, the formation energy of a given configuration is defined as:
$E_{\mathrm{f}}(x)=E(\mathrm{W_{x}Mo_{1-x}Te_{2}})-x\,E(\mathrm{\mathrm{WTe_{2}}})-(1-x)\,E(\mathrm{\mathrm{MoTe_{2}}}),$ (5)
where all reference energies are computed in the most stable polymorph at zero temperature (H for $\mathrm{MoTe_{2}}$ and $\mathrm{T^{\prime}}$ for $\mathrm{WTe_{2}}$). The quantity reported in Fig. 11 is the formation energy difference $E_{\mathrm{f}}^{\mathrm{H}}(x)-E_{\mathrm{f}}^{\mathrm{T^{\prime}}}(x)$, which directly determines the relative phase stability.
To generate validation configurations, we employ special quasirandom structures (SQS)[76] to approximate random W/Mo distributions on the metal sublattice at a given composition $x$. For each intermediate composition shown in Fig. 11, we construct and evaluate 10 independent SQS realizations to ensure statistical robustness of the estimated formation-energy differences.
As shown in Fig. 11, existing foundation models struggle with this task. MATTERSIM (v1.0.0–1M) and MACE (MATPES–PBE–0) incorrectly predict the $T^{\prime}$ phase to be the ground state for all compositions, while PET–MAD (v1.0.2), although capturing a phase transition and quite well the formation energy difference for end-members, places the critical concentration around $x\sim 0.7$—far from the DFT reference value at about $x\sim 0.3$.
Then, we employ AiiDA-TrainsPot to train three additional MACE models, all starting from the same four pristine input structures (H- and $\mathrm{T^{\prime}}$-phase $\mathrm{MoTe_{2}}$ and $\mathrm{WTe_{2}}$) taken from the Materials Cloud 2D database [43]:
  1. 1.
    Alloy model trained from scratch. A dataset of 1687 alloy configurations is generated by randomly sampling the W/Mo distribution in both phases during the dataset-augmentation stage, and a committee of models is trained from scratch on this dataset.
  2. 2.
    Fine-tuned alloy model. The same alloy dataset is used to fine-tune the MACE MATPES–PBE–0 foundation model.
  3. 3.
    End-member model with active learning. A fully independent model was trained from scratch using only configurations of the pure end-members in both phases (no alloys, clusters, surfaces, or substitutional configurations were generated during dataset augmentation). Twelve active-learning iterations are performed, yielding a final dataset of 6965 structures.
All three models markedly outperform the existing foundation models, in particulr accurately capturing both the composition dependence of the formation-energy difference and the H-T phase-stability crossover. Remarkably, the best-performing model is the one trained only on end-member configurations: despite never having seen alloy structures, it correctly reproduces the formation-energy trend and predicts the phase-stability crossover close to the DFT reference value. This result highlights the effectiveness of active learning with calibrated committee disagreement—and, more generally, of the AiiDA-TrainsPot protocol—in sampling the relevant regions of configuration space and constructing a compact dataset that yields a highly transferable NNIP.
Figure 11: Comparison of the formation-energy difference $E_{\mathrm{f}}^{\mathrm{H}}(x)-E_{\mathrm{f}}^{\mathrm{T^{\prime}}}(x)$ as a function of W concentration $x$ for several NNIPs, comparing foundation models with models trained or fine-tuned through AiiDA-TrainsPot. The MACE potentials generated with AiiDA-TrainsPot are trained on datasets produced via data augmentation starting from pure $\mathrm{MoTe_{2}}$ and $\mathrm{WTe_{2}}$ in both H and $\mathrm{T^{\prime}}$ phases. The models explicitly trained or fine-tuned on alloys share the same dataset of 1687 alloy structures generated in the dataset-augmentation stage. In contrast, the active-learning runs on end-members alloys only, i.e., no clusters, surfaces, nor substitutional configurations are generated during dataset augmentation and hence are not present in the training dataset. All DFT simulations are computed with the same level of theory, except for the training sets originally employed to build the foundation models by their respective authors.
III Conclusions
@@
The SOREP-based dimensionality reduction (tSNE) has shown that subsequent MD-based active learning steps explore the PES through a dual mechanism: dense sampling of already-known regions and, simultaneously, exploration of entirely new basins. A compelling example is the spontaneous formation of carbon nanotubes during the MD simulations in the active learning cycle: carbon nanotubes were absent from the original dataset but become then automatically incorporated by AiiDA-TrainsPot into subsequent training iterations, suggesting the ability of the automated workflow to find novel and quite different metastable or stable structures without prior knowledge. Indeed, the automated training strategy delivered potentials that were also able to reproduce the correct RDF of amorphous carbon.
@@
Furthermore, as demonstrated for both the carbon allotropes and the $\mathrm{W_{x}Mo_{1-x}Te_{2}}$ alloy benchmarks, AiiDA-TrainsPot can be effectively employed to fine-tune existing foundation models to specific applications. Fine-tuning can be performed either in a single shot on a fixed dataset or embedded into subsequent active-learning iterations, further improving model accuracy and robustness.
It would be interesting to investigate other data-augmentation approaches—such as random structure search [49], non-diagonal supercells [40], or generative models [41]—as well as advanced exploration strategies beyond NPT and NVT MD—such introducing metadynamics [38] by interfacing AiiDA-TrainsPot to PLUMED [16, 67, 62]. More in general, we hope that the AiiDA’s plugin system and the modular structure of AiiDA-TrainsPot will encourage and facilitate future upgrades, as well as the integration of new tools. An example would be supporting multiple NNIP backends (beyond MACE and Metatrain) and electronic structure codes (beyond Quantum ESPRESSO), in the spirit of previous efforts on code-agnostic common workflows for EOS and dissociation curves [32]. Powerful upgrades would be enabled by interfacing AiiDA-TrainsPot with existing specialized AiiDA workflows, for instance using DFT+U calculations where the Hubbard U can be automatically calculated for each configuration either with DFPT [64, 65, 66] through the AiiDA-Hubbard workflow [6] or even more efficiently through ML methods [69].
@@
While AiiDA-TrainsPot can operate autonomously for general-purpose NNIP development, domain experts retain full flexibility to incorporate their physical and chemical intuition in the automation strategy. The modular architecture is designed to enable full customization of all key components: initial structure selection, dataset augmentation parameters, MD simulation conditions, and computational settings for integrated codes—Quantum ESPRESSO, MACE, Metatrain and LAMMPS. Such level of customization allows users to balance the power of automation with the flexibility that is needed to support a wide range of applications, which require tailoring the active learning process to specific research objectives. Indeed, while AiiDA-TrainsPot automates the entire process—traditionally long, tedious and prone to human errors—of developing NNIPs, optimal results still benefit from careful consideration of the system of interest. In other words, the selection of initial structures, augmentation strategies, and MD conditions, can—and often should—be tailored to reflect the target application and desired properties: AiiDA-TrainsPot makes that effort straightforward. On top of that, the enforcement of standardized protocols (either already established, e.g., SSSP pseudopotentials [51] or “fast”/“stringent” QE protocols [45], or introduced in this work) contribute to precision, reproducibility and seamless integration with future efforts in training larger models.
@@
IV Methods
All calculations were performed using AiiDA-TrainsPot, Quantum ESPRESSO [25, 24, 26] v7.3.1, MACE [7] v0.3.12, metatrain [13] v2025.10, and LAMMPS [63] v8Feb2023.
Carbon allotropes
The initial dataset for carbon allotropes included 48 structures, primarily sourced from the Materials Cloud 2D and 3D databases [43, 31], with additional nanoribbons and fullerenes from the Drautz dataset [53].
@@ IV Methods
SACADA dataset
DFT calculations employed both the SSSP PBE and PBEsol precision pseudopotential libraries (v1.3) [51, 14, 20], using the same computational parameters as those adopted in the active learning. For fine-tuning, the PET-MAD potential (v1.0.2) [42] was trained using the default hyperparameters, with a batch size of 3 and 100 training epochs.
$\mathrm{W_{x}Mo_{1-x}Te_{2}}$ alloys
The initial dataset for $\mathrm{W_{x}Mo_{1-x}Te_{2}}$ alloys included 4 structures comprising the H and $\mathrm{T^{\prime}}$ polymorphs of pristine $\mathrm{MoTe_{2}}$ and $\mathrm{WTe_{2}}$ monolayers, sourced from the Materials Cloud 2D database [43].
During dataset augmentation, structures were replicated up to a minimum cell length of 18 Å. The non-periodic direction was padded with 15 Å of vacuum. Vacancies (2 per structure) were created in 30% of the structures. No substitutional configurations, clusters, or surfaces were generated during dataset augmentation. Only for the alloys dataset, random W/Mo distributions on the metal sublattice were sampled by setting in dataset augmentation stage Te as fixed_species and Mo and W as alloy_species.
The same DFT computational settings as in run B for carbon allotropes were used also in this case, while MD simulations performed in NPT ensemble (with barostat applied only to in-plane directions) explored temperatures ranging from 0 to 1800 K and pressures from 10-10 to $10$ kbar, without additional VdW corrections. MACE trainings were performed with radial cutoff of 6.0 Å, two message-passing layers, batch size of 5, and up to 500 epochs, except for fine-tuning run for which same parameter of the original MACE MATPES–PBE–0 model were used.
All the calculations in this work were performed on the CINECA Leonardo supercomputer, using nodes equipped with 4 NVIDIA A100 GPUs. Taking as reference the active-learning run used to train the MACE potential for MoTe2 and WTe2 (alloy benchmark in Fig. 11), where each structure contains about 108 atoms, we estimate an average computational cost of $\sim 1.2$ GPU hours per structure for DFT calculations, $\sim 25$ GPU hours for each MACE training run, and $\sim 0.1$ GPU hours for each LAMMPS MD simulation. After 12 active-learning iterations, resulting in a final dataset of 6965 structures, the total computational cost for training this committee of MACE potentials amounts to about 11 000 GPU hours, 84% of which was spent on DFT labelling.
Author Contributions
Conflicts of interests
There are no conflicts to declare.
Data availability
The training datasets and trained models are available on the Materials Cloud Archive [12]. AiiDA-TrainsPot is available on GitHub at https://github.com/aiida-trieste-developers/aiida-trains-pot.
Acknowledgements